# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ Abstract.
We study $\mathsf{RCD}$-spaces $(X,d,\mathfrak{m})$ with group actions by isometries preserving the reference measure $\mathfrak{m}$ and whose orbit space has dimension one, i.e. cohomogeneity one actions. To this end we prove a Slice Theorem asserting that each slice at a point is homeomorphic to a non-negatively curved $\mathsf{RCD}$-space. Under the assumption that $X$ is non-collapsed we further show that the slices are homeomorphic to metric cones over homogeneous spaces with $\operatorname{Ric}\geq 0$. As a consequence we obtain complete topological structural results and a principal orbit representation theorem. Conversely, we show how to construct new $\mathsf{RCD}$-spaces from a cohomogeneity one group diagram, giving a complete description of $\mathsf{RCD}$-spaces of cohomogeneity one. As an application of these results we obtain the classification of cohomogeneity one, non-collapsed $\mathsf{RCD}$-spaces of essential dimension at most $4$.
@@ 1. Introduction
The theory of $\mathsf{RCD}$-spaces has emerged from, and is still largely motivated, by the extensive and deep work of Cheeger and Colding on Gromov-Hausdorff limit spaces of sequences consisting of Riemannian manifolds with Ricci curvature bounded from below [18, 19, 20]. Cheeger and Colding themselves asked whether a synthetic treatment of Ricci curvature bounds was possible, the most natural answer being the now well developed theory of spaces with the curvature-dimension condition, initiated by Lott, Sturm and Villani [56, 74, 75] and refined to exclude Finslerian spaces by Ambrosio, Gigli and Savaré [2, 32]. Thus, the study of $\mathsf{RCD}$-spaces carries intrinsic theoretical value; however it has also produced several non-trivial applications to the study of Riemannian manifolds which are beyond the scope of the classical standpoint, i.e. only using smooth tools (see [33, Section 7] for some examples of these results).
@@
An approach to produce such a systematic method for the study of $\mathsf{RCD}$-spaces is the so called extended symmetry program. Grove originally proposed in [38] that in the search of new examples or obstructions to the existence of Riemannian metrics with positive sectional curvature we should first study the manifolds with a high degree of symmetry. This approach has been extended to the study of Riemannian manifolds with other lower curvature bounds, and has also been successfully extrapolated to the study of Alexandrov spaces, i.e. length spaces with a synthetic lower curvature bound (see [28, 44] and references therein).
@@
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space. Let $G$ be a compact Lie group acting on $X$ by measure preserving isometries and cohomogeneity one. Denote by $X^{\ast}$ the quotient space $\pi\colon X\to X/G$ equipped with the quotient metric $d^{\ast}(\pi(x),\pi(y)):=d(G(x),G(y))$. Then the following hold:
@@
    When $X^{\ast}$ is homeomorphic to $[0,1]$, then $X$ is equivariantly homeomorphic to the union of two cone bundles.
@@
    When $X^{\ast}$ is homeomorphic to $[0,\infty)$, then $X$ is equivariantly homeomorphic to a cone bundle.
@@
    When $X^{\ast}$ is homeomorphic to $\mathbb{R}$, then $X$ is equivariantly homeomorphic to ×$\mathbb{R}\times G/H$.
@@
If the tangent cones of $X$ at some points in the singular orbits $G(x_{+})$, $G(x_{-})$ have a metric cone structure, then the cone fibers are cones over $\mathsf{RCD}(N-k_{\pm}-2,N-k_{\pm}-1)$-spaces, where $k_{\pm}=\dim(G(x_{\pm}))$.
This result gives the full description of $\mathsf{RCD}$-spaces of cohomogeneity $1$, completing the rigidity results of Galaz-García, Kell, Mondino and Sosa in [27, Corollary 1.4]. Also, from the proof it is clear that in the case when we have two spaces $X_{1}$, $X_{2}$ with the same orbit space, and same orbit types, then they are equivariantly homeomorphic. We also show that to obtain the conclusions about the homeomorphism type in Theorem A we only need the geodesic space structure (see Section 3 and Theorem 4.12).
The topological description depends on having a geodesic structure, and the geometry of the quotient space, as stated in Theorem 4.12. Moreover, similar conclusions hold when $G$ is non-compact but acts properly by isometries, see Remark 4.13.
Moreover, in the case of Theorem A (a) we get three subgroups $H$, $G_{-}$, $G_{+}$, that up to conjugation are all the possible isotropy groups of the action. By Theorem 3.14 it follows that $H<G_{\pm}$, and thus we have a tuple $(G,H,G_{-},G_{+})$ associated to the action. We show that the converse is true: for each cohomogeneity one group diagram $(G,H,G_{-},G_{+})$, that is a collection of subgroups where $H$ is a Lie subgroup of $K_{\pm}$, and in turn $K_{\pm}$ are Lie subgroups of $G$ and such that $K_{\pm}/H$ are homogeneous spaces with positive Ricci curvature, determines a cohomogeneity one $\mathsf{RCD}$-space.
@@
Let $G$ be a compact Lie group and $K_{+}$, $K_{-}$, $H$ Lie subgroups of $G$ such that $(G,H,K_{+},K_{-})$ is a cohomogeneity one group diagram. Then, there exists an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ admitting a cohomogeneity one action of $G$ by measure preserving isometries such that the associated group diagram is $(G,H,K_{+},K_{-})$.
These two results give a complete description of $\mathsf{RCD}$-spaces of cohomogenity $1$, which is analogous to the descriptions of topological and Riemannian manifolds of cohomogeneity $1$, and Alexandrov spaces of cohomogeneity $1$ (see [61], [60], [29], and [28]).
A crucial ingredient in the proof of the classification Theorem A is the Slice Theorem. This is used in general in the Grove symmetry program as a recognition tool for the local topology around a fixed orbit.
All known proofs of the Slice Theorem, be it in the smooth or Alexandrov setting, require a notion of “space of directions” and exponential map (see [44]). In the $\mathsf{RCD}$-setting, while a notion of space of directions can be defined a.e. (see [42]), the infinitesimal-to-local relation (in the spirit of Perelman’s conical neighborhood theorem) is a priori not known to hold. That is, it is not yet fully understood how the mGH-tangent space and $L^{2}$-tangent module (which are equivalent by work of Gigli and Pasqualetto [37]) determine the space locally (see however the recent work of Honda and Peng [49] in this direction). We instead work in the setting of optimal transport theory, namely using ideas from needle-decomposition [15], and also exploit the low-dimensionality of the orbit space to circumvent this obstacle.
@@
Let $(X,d)$ be a geodesic space and $G$ a compact Lie group acting on $X$ by isometries such that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then for any $x_{0}\in X$, there exists $\delta_{0}>0$ such that for any $0<\delta<\delta_{0}$ we have $B_{\delta}(x_{0})\subset\mathcal{T}^{nb}_{u}$ and
$S_{x_{0}}:=\left(\Gamma^{-}_{u}(x_{0})\cap B_{\delta}(x_{0})\right)\cup\{x_{0}\}$
is a slice through $x_{0}$.
Here we used the customary definitions as follows. Denoting $u(x):=d(x,G(x_{0}))$, we set $\Gamma^{-}_{u}(x)=\{y\in X\mid(x,y)\in\Gamma^{-}_{u}\}$, where as usual, $\Gamma^{-}_{u}=\left\{(x,y)\mid u(y)-u(x)=d(x,y)\right\}$. We also denote the so-called non-branching transport set by $\mathcal{T}^{nb}_{u}$. It is worth pointing out that we do not assume the $\mathsf{RCD}$-condition (or any other bounded curvature condition) in this result and thus in particular it does not depend on any specific bound of the curvature. It depends only on the dimension of the orbit space and properties of the metric structure.
Once this description of the slices at each point has been established we move to show that in general the slices admit a non-negatively curved $\mathsf{RCD}$-structure (that is, each slice is homeomorphic to an $\mathsf{RCD}(0,M)$-space, for the appropriate dimension bound $M$ as stated below). When specializing to the case where all tangent cones of $X$ are metric cones we further show that slices are homeomorphic to metric cones over homogeneous spaces with adequate $\mathsf{RCD}$-structures. This condition is satisfied for example when $X$ is non-collapsed [23, Proposition 2.8].
@@
Let $(X,d,\mathfrak{m}_{X})$ be an $\mathsf{RCD}(K,N)$-space with $N>2$, such that all tangent cones are metric cones. Assume that $G$ is a compact Lie group acting effectively by measure preserving isometries on $X$, such that $(X^{\ast},d^{\ast},\mathfrak{m}_{X}^{\ast})$ has essential dimension equal to $1$. Then the following hold:
@@
    For any $x_{0}\in X$ the slice $S_{x_{0}}$ admits an $\mathsf{RCD}(0,N-k)$ structure, where $k=\dim(G(x_{0}))$.
@@
    Moreover, $S_{x_{0}}$ is homeomorphic to a metric cone over a homogeneous smooth Riemannian manifold with Ricci curvature greater or equal to $N-k-2$.
We point out that by the work of Palais [65] the existence of slices is well understood for compact group actions by isometries. Nonetheless, in the Riemannian and Alexandrov setting the metric structure determines a canonical choice of a slice, which inherits geometric properties of the global space. In this way, for the local topological recognition we may apply a reduction on the dimension being considered while still considering spaces in the appropriate setting. This key point is beyond the conclusions of Palais’ Slice Theorem [65], and it is the point being proven in Theorem C and Theorem D. Thus, by Theorem D, Theorem C gives a positive answer to [22, Problem 1] for non-collapsed $\mathsf{RCD}$-spaces of cohomogeneity one.
The slice in Theorem D is the fiber of the cone bundles in Theorem A. In the non-collapsed case, the homogeneous spaces are determined by the slice.
Finally, we use our previous results to obtain a topological classification of non-collapsed $\mathsf{RCD}$-spaces of essential dimensions at most $4$, in the spirit of the corresponding classification in the smooth and Alexandrov cases. In fact, we obtain that all the $\mathsf{RCD}$-spaces considered with these restrictions are homeomorphic to Alexandrov spaces. This solves a particular case (that of spaces admitting cohomogeneity one actions) of a conjecture by Mondino asserting that $\mathsf{RCD}$-spaces with essential dimension $3$ are homeomorphic to orbifolds (see Question 9 in Section 2 of [22]).
@@
Let $G$ be a compact Lie group acting almost effectively by measure preserving isometries and cohomogeneity one on a closed non-collapsed $\mathsf{RCD}(K,N)$-space $(X,d,\mathfrak{m})$ with $N\leq 4$. Then $X$ is homeomorphic to an Alexandrov space.
Combining the characterization in [6] of homogeneous spaces admitting positive Ricci curvature with the cone construction in [51]. it is easy to construct $\mathsf{RCD}(K,N)$-spaces of cohomogeneity one for some $K\geq 0$. This construction allow us to show that that the dimension bound in Theorem E is optimal (see Theorem 7.1). Namely the following example shows that in dimension $5$ there are non-collapsed $\mathsf{RCD}$-spaces of cohomogeneity one which are not Alexandrov spaces of cohomogeneity one:
@@
The suspension $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits an non-collapsed $\mathsf{RCD}(K,5)$-structure for any $K\geq 0$ such that the suspension of the $\mathrm{SO}(3)\times\mathrm{SO}(3)$-action on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is by measure preserving isometries and by cohomogeneity one, with group diagram $(\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(2)\times\mathrm{SO}(2),\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(3)\times\mathrm{SO}(3))$. Moreover, $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ with this action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is not an Alexandrov space of cohomogeneity one, i.e. it does not admit a metric making it an Alexandrov space and such that the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is by isometries.
We observe that in the previous example we obtain $\mathsf{RCD}$-spaces of cohomogeneity one that are not non-collapsed, and the same holds true for spherical suspensions of homogeneous spaces with sufficient curvature, see Theorem 7.1. Nonetheless, we observe that since these examples are $\mathsf{RCD}$-spaces of cohomogeneity one, the stronger conclusions of Theorem D hold by construction. Thus we ask the following question:
@@
Does there exist an $\mathsf{RCD}$-space $X$ equipped with a measure preserving isometric action of cohomogeneity one by a compact Lie group $G$, which is not non-collapsed, and such that the slice at any orbit does not admit an metric cone structure over a homogeneous space?
In other words, given an $\mathsf{RCD}$-space $X$ equipped with a measure preserving isometric action of cohomogeneity one by a compact Lie group $G$, does it always admit an non-collapsed invariant $\mathsf{RCD}$-structure?
We also observe that the spaces in Theorem A (c), and (d) admit an $\mathsf{RCD}(0,N)$-structure of cohomogeneity one, given by a smooth Riemannian metric on the smooth representatives of the homoeomrphism types. Moreover, due to Gigli’s Splitting Theorem [31] the spaces in Theorem A (d) do not admit an $\mathsf{RCD}(K,N)$-structure with $K>0$. For spaces in Theorem A (c), since they have infinite fundamental group, by Myers Theorem (see [59, Theorem 3.5] toghether with [78, Main Theorem]) it follows that they also do not admit an $\mathsf{RCD}(K,N)$-structure with $K>0$. For spaces in Theorem A (b), we give in Theorem I sufficient conditions to guarantee that they also admit an $\mathsf{RCD}(0,N)$-structure of cohomogeneity one. Thus it natural to ask if this also holds for the remaining cases:
@@
Let $X$ be an $\mathsf{RCD}$-space as in Theorem A (a) or (b). Does it admit a (possibly different) new $\mathsf{RCD}(0,N)$-structure of cohomogeneity one?
This question was answered in the positive by Grove and Ziller in [39, Theorem B] when $X$ is smooth Riemannian manifold. Moreover, in [39, Theorem A], they showed that a smooth compact Riemannian manifold with an action of cohomogeneity one admits a Riemannian metric of positive Ricci curvature if and only if it has finite fundamental group. Due to the fact that Myers Theorem holds for $\mathsf{RCD}$-spaces, we ask if the cardinality of the fundamental group is the unique obstruction to “synthetic positive Ricci curvature” in the case of compact $\mathsf{RCD}$-spaces of cohomogeneity one.
Question 1.3.
Let $X$ be an $\mathsf{RCD}$-space as in Theorem A (a). If $X$ has finite fundamental group, does it admit an $\mathsf{RCD}(K,N)$-structure with $K>0$?
We present a simple presentation of the fundamental group of an $\mathsf{RCD}$-space for the form Theorem A (a) in Theorem 4.14.
The organization of the article is the following. In Section 2 we collect the definitions and results from the theory of non-smooth differential geometry and transformation groups that we will need in the sequel. In Section 3 we prove Theorem C. We also prove two results showing that our construction using the transport set is a slice without the cohomogeneity one restriction, albeit in the case that the orbit space happens to be isometric to an Alexandrov space, which are of independent interest. In Section 4 we prove Theorem D obtaining along the way a version of the orbit representation theorem. We also establish the topological rigidity result contained in Theorem A. We continue in Section 5 proving that one can construct $\mathsf{RCD}$-spaces from any given cohomogeneity one group diagram, as expressed in Theorem B. Finally, we address the low-dimensional classification of Theorem E in Section 6.
Acknowledgements The authors wish to thank Fernando Galaz-García, Nicola Gigli, John Harvey, Alexander Lytchak, Marco Radeschi and Masoumeh Zarei for very useful discussions and communications and for commenting on a first draft of this manuscript. We thank Elia Bruè for useful discussions about the infinitesimal action which improved section 4, and for pointing us to the work of Pan and Wei [66].

### v3

@@ Abstract.
We study $\mathsf{RCD}$-spaces $(X,d,\mathfrak{m})$ with group actions by isometries preserving the reference measure $\mathfrak{m}$ and whose orbit space has dimension one, i.e. cohomogeneity one actions. To this end we prove a Slice Theorem asserting that when $X$ is non-collapsed the slices are homeomorphic to metric cones over homogeneous spaces with $\operatorname{Ric}\geq 0$. As a consequence we obtain complete topological structural results (also in the collapsed case) and a regular orbit representation theorem. Conversely, we show how to construct new $\mathsf{RCD}$-spaces from a cohomogeneity one group diagram, giving a complete description of $\mathsf{RCD}$-spaces of cohomogeneity one. As an application of these results we obtain the classification of cohomogeneity one, non-collapsed $\mathsf{RCD}$-spaces of essential dimension at most $4$.
@@ 1. Introduction
The theory of $\mathsf{RCD}$-spaces has emerged from, and is still largely motivated, by the extensive and deep work of Cheeger and Colding on Gromov-Hausdorff limit spaces of sequences consisting of Riemannian manifolds with Ricci curvature bounded from below [17, 18, 19]. Cheeger and Colding themselves asked whether a synthetic treatment of Ricci curvature bounds was possible. The most natural answer is the now well-developed theory of spaces with the curvature-dimension condition, initiated by Lott, Sturm, and Villani [57, 78, 79] and refined to exclude Finslerian spaces by Ambrosio, Gigli, and Savaré [5, 30]. Thus, the study of $\mathsf{RCD}$-spaces carries intrinsic theoretical value; however, it has also produced several non-trivial applications to the study of Riemannian manifolds which are beyond the scope of the classical standpoint, i.e. only using smooth tools (see [31, Section 7] for some examples of these results).
@@
An approach to produce such a systematic method for the study of $\mathsf{RCD}$-spaces is the so called extended symmetry program. Grove originally proposed in [36] that, in the search of new examples or obstructions to the existence of Riemannian metrics with positive sectional curvature, we should first study the manifolds with a high degree of symmetry. This approach has been extended to the study of Riemannian manifolds with other lower curvature bounds, and has also been successfully extrapolated to the study of Alexandrov spaces, i.e. length spaces with a synthetic lower sectional curvature bound (see [26, 42] and references therein).
@@
    When $X^{\ast}$ is homeomorphic to $[0,1]$, then $X$ is equivariantly homeomorphic to the union of two cone bundles over the singular orbits, glued along their boundary.
@@
    When $X^{\ast}$ is homeomorphic to $0,\infty)$, then $X$ is homeomorphic to a cone bundle.
@@
    When $X^{\ast}$ is homeomorphic to $\mathbb{R}$, then $X$ is homeomorphic to $\mathbb{R}\times G/H$.
@@
In the case when $X$ is non-collapsed then the cone fibers admit metric cone structures over $\mathsf{RCD}(N-k_{\pm}-2,N-k_{\pm}-1)$-spaces, where $k_{\pm}=\dim(G(x_{\pm}))$.
This result gives the full description of $\mathsf{RCD}$-spaces of cohomogeneity $1$, completing the rigidity results of Galaz-García, Kell, Mondino and Sosa in [25, Corollary 1.4]. Also, it will be clear from the proof that if $X_{1}$ and $X_{2}$ have the same orbit space and orbit types then they are equivariantly homeomorphic. We also show that, to obtain the conclusions about the homeomorphism type in ?THM? LABEL:MC:_Homeomorphism_rigidity, we only need an essentially non-branching metric-measure structure (see Section 3 and ?THM? LABEL:T:_Topological_classification_geodesic_coho_1_spaces.).
The topological description depends on having a geodesic structure, and the geometry of the quotient space, as stated in ?THM? LABEL:T:_Topological_classification_geodesic_coho_1_spaces..
Moreover, in the case of ?THM? LABEL:MC:_Homeomorphism_rigidity (a) we get three subgroups $H$, $K_{-}$, and $K_{+}$, that, up to conjugation, are all the possible isotropy groups of the action. By ?THM? LABEL:T:_Principal_isotropy_is_subgroup_of_singular_isotropy, it follows that $H<K_{\pm}$. Thus, we have a tuple $(G,H,K_{-},K_{+})$ associated to the action. We show that the converse is true: each cohomogeneity one group diagram $(G,H,K_{-},K_{+})$, that is a collection of subgroups where $H$ is a Lie subgroup of $K_{\pm}$, and in turn $K_{\pm}$ are Lie subgroups of $G$ such that $K_{\pm}/H$ are homogeneous spaces with positive Ricci curvature if they have dimension $\geq 2$, and in the case of dimension $1$ they have bounded diameter, determines a cohomogeneity one $\mathsf{RCD}$-space.
@@
Let $G$ be a compact Lie group and $_{+}, KK_{-}$, $H$ Lie subgroups of $G$ such that $(G,H,K_{+},K_{-})$ is a cohomogeneity one group diagram. Then, there exists an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ admitting a cohomogeneity one action of $G$ by measure preserving isometries such that the associated group diagram is $(G,H,K_{+},K_{-})$.
These two results give a complete description of $\mathsf{RCD}$-spaces of cohomogeneity $1$, which is analogous to the descriptions of topological and Riemannian manifolds of cohomogeneity $1$, and Alexandrov spaces of cohomogeneity $1$ (see [1], [63], [62], [66], [27], and [26]).
We also remark that the condition on $G_{\pm}/H$ having a metric of positive Ricci curvature is equivalent to $G_{\pm}/H$ being compact and with finite fundamental group [9, Theorem 1], in the case of dimensions at least $2$.
A crucial ingredient in the proof of the classification ?THM? LABEL:MC:_Homeomorphism_rigidity is the Slice Theorem. This is used in general in the Grove symmetry program as a recognition tool for the local topology around giving an explicit topological model of a tubular neighborhood of a fixed orbit.
All known geometric proofs of the Slice Theorem, be it in the smooth or Alexandrov setting, require a notion of “space of directions” and exponential map (see [42]). In the $\mathsf{RCD}$-setting, while a notion of space of directions can be defined a.e. (see [40]), the infinitesimal-to-local relation (in the spirit of Perelman’s conical neighborhood theorem) is a priori not known to hold. That is, it is not yet fully understood how the mGH-tangent space and $L^{2}$-tangent module (which are equivalent by work of Gigli and Pasqualetto [35]) determine the space locally (see however the recent work of Honda and Peng [48] in this direction). Instead, we work in the setting of optimal transport theory and exploiting the space’s low-dimensionality to circumvent this obstacle.
@@
Let $(X,d,\mathfrak{m})$ be an essentially non-branching space (see ?THM? LABEL:D:_essentiallynonbranch ) and $G$ a compact Lie group acting on $X$ by measure-preserving isometries such that $X^{\ast}$ is isometric to [1,1][-1,1], $0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then for any $x_{0}\in X$, there exists $\delta_{0}>0$ such that, for any $0<\delta<\delta_{0}$ we have that
$S_{x_{0}}:=\{y\in X\,|\,d(x_{0},y)=d(G(x_{0}),y)\}\cap B_{\delta}(x_{0})$
is a slice through $x_{0}$ (see ?THM? LABEL:D:_Slice).
It is worth pointing out that we do not assume the $\mathsf{RCD}$-condition (or any other bounded curvature condition) in this result. Thus, it does not depend on any specific curvature bound. It depends only on the dimension of the orbit space and properties of the metric measure structure.
Once this description of the slices at each point has been established we move on to show that, in general, the slices admit a non-negatively curved $\mathsf{RCD}$-structure (that is, each slice is homeomorphic to an $\mathsf{RCD}(0,M)$-space, for the appropriate dimension bound $M$ as stated below). When considering the case when $X$ is non-collapsed, we further show that slices are homeomorphic to metric cones over homogeneous spaces with adequate $\mathsf{RCD}$-structures.
@@
Let $(X,d,\mathcal{H}^{N})$ be a non-collapsed $\mathsf{RCD}(K,N)$-space with $N>2$. Assume that $G$ is a compact Lie group acting effectively by measure preserving isometries on $X$, such that $(X^{\ast},d^{\ast},(\mathcal{H}^{N})^{\ast})$ has essential dimension equal to $1$. Then the following hold:
@@
    For any $x_{0}\in X$, the slice $S_{x_{0}}$ admits an $\mathsf{RCD}(0,N-k)$ structure, where $k=\dim(G(x_{0}))$.
@@
    Moreover, $S_{x_{0}}$ is homeomorphic to a metric cone over a homogeneous smooth Riemannian manifold with Ricci curvature greater or equal to Nk2N-k-2.
We point out that by the work of Palais [68] the existence of slices is well understood for compact group actions by isometries. Nonetheless, in the Riemannian and Alexandrov setting, the metric structure determines a canonical choice of a slice, which inherits geometric properties of the global space. In this way, for the local topological recognition, we may apply a reduction on the dimension being considered while still considering spaces in the appropriate setting. This key point is beyond the conclusion of Palais’ Slice Theorem [68], and it is the point being proven in ?THM? LABEL:MT:_Slice_Theorem_Cohomogenity_one and ?THM? LABEL:MT:_Geometry_of_the_slice. Thus, by ?THM? LABEL:MT:_Geometry_of_the_slice, ?THM? LABEL:MT:_Slice_Theorem_Cohomogenity_one gives a positive answer to [21, Problem 1] for non-collapsed $\mathsf{RCD}$-spaces of cohomogeneity one.
The slice in ?THM? LABEL:MT:_Geometry_of_the_slice is the fiber of the cone bundles in ?THM? LABEL:MC:_Homeomorphism_rigidity. In the non-collapsed case, the homogeneous spaces are determined by the slice.
Finally, we use our previous results to obtain a topological classification of non-collapsed $\mathsf{RCD}$-spaces of essential dimensions at most $4$, in the spirit of the corresponding classifications in the smooth and Alexandrov cases. In fact, we obtain that all the $\mathsf{RCD}$-spaces considered with these restrictions are homeomorphic to Alexandrov spaces. This solves a particular case (that of spaces admitting cohomogeneity one actions) of a conjecture by Mondino asserting that $\mathsf{RCD}$-spaces with essential dimension $3$ are homeomorphic to orbifolds (see Question 9 in Section 2 of [21]).
@@
Let $G$ be a compact Lie group acting almost effectively by measure preserving isometries and cohomogeneity one on a closed non-collapsed $\mathsf{RCD}(K,N)$-space $(X,d,\mathcal{H}^{N})$ with $N\leq 4$. Then $X$ is homeomorphic to an Alexandrov space.
Combining the characterization in [9] of homogeneous spaces admitting positive Ricci curvature with the cone construction in [50], it is easy to construct $\mathsf{RCD}(K,N)$-spaces of cohomogeneity one for some $K\geq 0$. This construction allows us to show that the dimension bound in ?THM? LABEL:MT:_Low-dim-classification is optimal (see ?THM? LABEL:T:_Suspensions_are_RCD_spaces_of_coho_1). Namely, the following example shows that in dimension $5$ there are non-collapsed $\mathsf{RCD}$-spaces of cohomogeneity one which are not Alexandrov spaces of cohomogeneity one.
@@
The suspension $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits a non-collapsed $\mathsf{RCD}(K,5)$-structure for any $K\geq 0$ such that the suspension of the $\mathrm{SO}(3)\times\mathrm{SO}(3)$-action on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is by measure preserving isometries and by cohomogeneity one, with group diagram $(\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(2)\times\mathrm{SO}(2),\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(3)\times\mathrm{SO}(3))$. Moreover, $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ with this action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is not an Alexandrov space of cohomogeneity one, i.e. it does not admit a metric making it an Alexandrov space and such that the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is by isometries.
We observe that the spaces in ?THM? LABEL:MC:_Homeomorphism_rigidity (c) and (d) admit an $\mathsf{RCD}(0,N)$-structure of cohomogeneity one given by a smooth Riemannian metric on the smooth representatives of the homeomorphism types. Moreover, due to Gigli’s Splitting Theorem [29], the spaces in ?THM? LABEL:MC:_Homeomorphism_rigidity (d) do not admit an $\mathsf{RCD}(K,N)$-structure with $K>0$. For spaces in ?THM? LABEL:MC:_Homeomorphism_rigidity (c), since they have infinite fundamental group, by Myers Theorem (see [61, Theorem 3.5] together with [81, Main Theorem]) it follows that they also do not admit an $\mathsf{RCD}(K,N)$-structure with $K>0$. For spaces in ?THM? LABEL:MC:_Homeomorphism_rigidity (b), we give in Theorem LABEL:MT:_coho_one_ray_with_conditions_is_RCD_0,N sufficient conditions to guarantee that they also admit an $\mathsf{RCD}(0,N)$-structure of cohomogeneity one. Thus it natural to ask if this also holds for the remaining cases.
@@
Let $X$ be an $\mathsf{RCD}$-space as in ?THM? LABEL:MC:_Homeomorphism_rigidity (a) or (b). Does it admit a (possibly different) $\mathsf{RCD}(0,N)$-structure of cohomogeneity one?
This question was answered positively by Grove and Ziller in [37, Theorem B] when $X$ is a smooth Riemannian manifold. Moreover, in [37, Theorem A], they showed that a smooth compact Riemannian manifold with an action of cohomogeneity one admits a Riemannian metric of positive Ricci curvature if and only if it has finite fundamental group. Due to the fact that Myers Theorem holds for $\mathsf{RCD}$-spaces, we ask if the cardinality of the fundamental group is the unique obstruction to “synthetic positive Ricci curvature” in the case of compact $\mathsf{RCD}$-spaces of cohomogeneity one.
@@
Let $X$ be an $\mathsf{RCD}$-space as in ?THM? LABEL:MC:_Homeomorphism_rigidity (a). If $X$ has finite fundamental group, does it admit an $\mathsf{RCD}(K,N)$-structure with $K>0$?
We present a simple presentation of the fundamental group of an $\mathsf{RCD}$-space satisfying ?THM? LABEL:MC:_Homeomorphism_rigidity (a) in ?THM? LABEL:T:_fundamental_group.
The organization of the article is the following. In Section 2 we collect the definitions and results from the theory of non-smooth differential geometry and transformation groups that we will need in the sequel. In Section 3 we prove ?THM? LABEL:MT:_Slice_Theorem_Cohomogenity_one. We are also able to prove that the set $S_{x}$ is a slice without the cohomogeneity one restriction, albeit in the case that the orbit space happens to be isometric to an Alexandrov space, which is of independent interest. In Section 4 we prove ?THM? LABEL:MT:_Geometry_of_the_slice obtaining along the way a version of the orbit representation theorem. We also establish the topological rigidity result contained in ?THM? LABEL:MC:_Homeomorphism_rigidity. We continue in Section 5 proving that one can construct $\mathsf{RCD}$-spaces from any given cohomogeneity one group diagram, as expressed in ?THM? LABEL:thm.rcd_space_from_group_diagram. Finally, we address the low-dimensional classification of ?THM? LABEL:MT:_Low-dim-classification in Section 6.
Acknowledgments
The authors wish to thank Fernando Galaz-García, Nicola Gigli, John Harvey, Alexander Lytchak, Marco Radeschi, and Masoumeh Zarei for very useful discussions and communications and for commenting on a first draft of this manuscript. We thank Elia Bruè for useful discussions about the infinitesimal action which improved Section 4, and for pointing us to the work of Pan and Wei [69]. We are also grateful to Jiayin Pan and Dimitri Navarro for useful discussions on results in Section 4. We thank Christoph Böhm for questions about Section 5.